If you were moved 35ml of your solution and dilute the 35ml with water to a new volume of 250ml what is the new molarity
1
Expert's answer
2013-03-19T10:27:55-0400
In chemistry, the molar concentration,C is defined as the amount of a constituent n divided by the volume of the mixture V. The molarity is some amount per one liter.
If starting molarity is X M, and volume is 250 ml, the amount of compound is X*0.25 =0.25X mol.
The amount of compound in 35 ml is 0.035 L * X =0.035X. New amount (after adding) is 0.25X-0.035X =0,215X mol of compounds is still in the solution, even after adding new water.
So it is 0.215X mol in 250 ml, and it is 0.215X * 1000ml/250ml = 0.86X mol per one liter.
Molarity is 0.86X M